OpenStudy (anonymous):

Let f be a linear function, and let g be a quadratic function. The function f + g is _____. A. linear B. quadratic C. cubic D. None of the above

OpenStudy (anonymous):

Try out an example. f(x)=x+1 is a linear function, g(x)=x^2+x is a quadratic function, add them together and see what you get.

OpenStudy (anonymous):

its still a quadratic formula isent it?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

THANKS! I LOVE UR PANDA!

OpenStudy (anonymous):

LOL